# Kernel Ridge Riesz Representers: Generalization Error and Mis-specification

### v3

@@ 4.3 Standalone closed form solution
In summary, the standalone closed form for KRRR is not necessary for causal estimation in the special case of Corollary 4.4. However, Corollary 4.5 shows that if $\hat{\gamma}\neq\tilde{\gamma}$, or if $\hat{\gamma}$ or $\tilde{\alpha}$ is estimated on a held out sample, then a standalone closed form for KRRR that can be evaluated away from the training data is necessary for causal estimation. I derive it below using techniques of [Chernozhukov et al., 2020], as a secondary contribution.
Even for the special case of Corollary 4.4, this paper’s primary contribution is the population $L_{2}$ rate $\|\tilde{\alpha}-\alpha_{H}\|_{2}$ for kernel ridge balancing weights (and more generally KRRR) in Section 5. This generalization error is essential in order to appeal to strong semiparametric guarantees that allow some mis-specification and inference on causal functions.
In what follows, I use some additional notation for features. Define the operator $\Phi:H\rightarrow\mathbb{R}^{n}$ with $i$th component $\langle\phi(W_{i}),\cdot\rangle_{H}$ and likewise the operator $\Phi^{(m)}:H\rightarrow\mathbb{R}^{n}$ with $i$th component $\langle\phi^{(m)}(W_{i}),\cdot\rangle_{H}$. Concatenate $\Phi$ and $\Phi^{(m)}$ as $\Psi:H\rightarrow\mathbb{R}^{2n}$, with adjoint $\Psi^{*}:\mathbb{R}^{2n}\rightarrow H$.
Lemma 4.3 (Standalone closed form exists; c.f. Lemma 2 of [Chernozhukov et al., 2020]).
Suppose the conditions of Lemma 4.1 hold. There exists some $\rho\in\mathbb{R}^{2n}$ such that $\tilde{\alpha}=\Psi^{*}\rho$.
@@
The familiar kernel matrix $K^{(1)}=\Phi\Phi^{*}\in\mathbb{R}^{n\times n}$ encodes inner products between actual observations. The extended kernel matrix $K=\Psi\Psi^{*}\in\mathbb{R}^{2n\times 2n}$ encodes inner products between actual observations and counterfactual observations. Each entry of $K$ can be computed from data using the kernel $k$ and formula $m$ as discussed in Section 3.3. For intuition, write
@@
Proposition 4.1 (Standalone closed form for KRRR).
Suppose the conditions of Lemma 4.1 hold. Given observations $(W_{i})$ and evaluation location $w$, as well as the formula $m$, kernel $k$, and regularization $\lambda$, the KRRR closed form is as follows.
@@
    Calculate $\Omega^{2n\times 2n}$ and $u,v\in\mathbb{R}^{2n}$ by
@@
The standalone closed form solution for kernel ridge balancing weights, and KRRR more generally, appears to have been previously unknown. Proposition 4.1 demonstrates that it can be easily computed using only the formula $m$ and kernel $k$, without directly evaluating the feature maps $\phi(w)$ or $\phi^{(m)}(w)$. In this sense, it is a practical contribution for inference. See Appendix E for its specialization to Example 3.2.
5 Primary contribution: Generalization error with mis-specification
@@ 5.1 Main result
I now provide the main result: population $L_{2}$ rates for KRRR that coincide with the optimal rate, where optimal rates are known. I study the class of Riesz representers defined by Assumption 3.1, imposing familiar smoothness and spectral decay conditions for the RKHS setting defined by Assumptions 3.2 and 3.3. I also impose mild regularity conditions that help in the algorithm derivation, defined by Assumptions 4.1 and 4.2.
@@
Theorem 5.1 (Upper bound for KRRR).
Suppose Assumptions 3.13.23.3(i), 4.1, and 4.2 hold with $\bar{M}$ bounded above. Set
$\lambda=\begin{cases}n^{-\frac{1}{2}}&\text{ if }\quad b=\infty;\\ n^{-\frac{b}{bc+1}}&\text{ if }\quad b\in(1,\infty),\;c\in(1,2];\\ \ln^{\frac{b}{b+1}}(n)\cdot n^{-\frac{b}{b+1}}&\text{ if }\quad b\in(1,\infty),\;c=1.\end{cases}$
Then $\lim_{\tau\rightarrow\infty}\lim\sup_{n\rightarrow\infty}\sup_{p\in\mathcal{P}(b,c)}\mathbb{P}_{(W_{i})\sim p^{n}}\{\mathcal{R}(\tilde{\alpha})>\tau\cdot r_{n}\}=0$, where
$r_{n}=\begin{cases}n^{-1}&\text{ if }\quad b=\infty;\\ n^{-\frac{bc}{bc+1}}&\text{ if }\quad b\in(1,\infty),\;c\in(1,2];\\ \ln^{\frac{b}{b+1}}(n)\cdot n^{-\frac{b}{b+1}}&\text{ if }\quad b\in(1,\infty),\;c=1.\end{cases}$
Remark 5.1 (Causal functions).
When the causal parameter has a formula of the form $m_{h}(w,f)=\ell_{h}(w)\check{m}(w,f)$, as in Example 3.2, $\bar{M}_{h}\uparrow\infty$. In such case, I use the estimator $\hat{\alpha}_{h}(W)=\ell_{h}(W)\tilde{\alpha}(W)$, where $\tilde{\alpha}$ is KRRR for the formula $\check{m}(w,f)$. When $\tilde{\alpha}$ has the rate $r_{n}$, $\hat{\alpha}_{h}$ has the rate $h^{-2}r_{n}$ under weak regularity conditions. See Appendix A for details.
@@
Theorem 5.1 does not require Assumption 3.3(ii). Moreover, as an intermediate step, I prove a nonasymptotic bound (Proposition 6.2) without imposing Assumptions 3.2 and 3.3(i), keeping all constants explicit.
When the RKHS is finite dimensional, i.e. $b=\infty$ in Assumption 3.3, KRRR achieves the parametric rate $r_{n}=n^{-1}$. When the RKHS is infinite dimensional with at least polynomial spectral decay, i.e. $b<\infty$ in Assumption 3.3, KRRR achieves a familiar rate $r_{n}=n^{-\frac{bc}{bc+1}}$. This rate depends on $c$ in Assumption 3.2, which quantifies the smoothness of the best kernel approximation $\alpha_{H}\in H$ to the Riesz representer $\alpha_{0}\in L_{2}$. If no additional smoothness of $\alpha_{H}$ is known, i.e. all we have is that $\alpha_{H}\in H$ by construction, then $c=1$ and the rate has an extra logarithmic factor.
The rate $r_{n}=n^{-\frac{bc}{bc+1}}$ generalizes the familiar rate from Sobolev analysis. For example, take $\mathcal{W}=\mathbb{R}^{p}$ and denote by $\mathbb{H}_{2}^{s}$ the Sobolev space with $s>p/2$ square integrable derivatives, as before. Suppose that KRRR $\tilde{\alpha}$ is estimated using the RKHS $H=\mathbb{H}_{2}^{s}$. Suppose that $\alpha_{H}\in\mathbb{H}_{2}^{s_{0}}$, i.e. the best kernel approximation to the Riesz representer $\alpha_{0}\in L_{2}$ has $s_{0}>s$ square integrable derivatives. The population $L_{2}$ rate for $\tilde{\alpha}$ is then $r_{n}=n^{-\frac{2s_{0}}{2s_{0}+p}}$. When $s_{0}=s$, which is a lower bound on $s_{0}$ by construction, $r_{n}=\ln(n)^{\frac{2s}{2s+p}}\cdot n^{-\frac{2s}{2s+p}}$. Theorem 5.1 applies to settings beyond Sobolev spaces over Euclidean domains.
Corollary 5.1 (Special case: Kernel ridge regression; Theorem 1 of [Caponnetto and De Vito, 2007]).
Suppose the conditions of Theorem 5.1 hold, replacing Assumption 3.1 with the assumption that $|Y|\leq\bar{Y}$ almost surely. Then
$\lim_{\tau\rightarrow\infty}\lim\sup_{n\rightarrow\infty}\sup_{p\in\mathcal{P}(b,c)}\mathbb{P}_{(W_{i})\sim p^{n}}\{\mathcal{R}(\tilde{\gamma})>\tau\cdot r_{n}\}=0.$
Remark 5.2 (Excess risk).
@@
Just as KRRR generalizes kernel ridge regression, Theorem 5.1 generalizes [Caponnetto and De Vito, 2007, Theorem 1]. The contribution is meaningful because KRRR also generalizes the classical kernel ridge balancing weights, for which generalization error was not fully articulated. In settings where lower bounds are known, the upper bounds in Theorem 5.1 often achieve them.
Lemma 5.1 (Lower bound; Theorem 2 of [Caponnetto and De Vito, 2007]).
Suppose the conditions of Corollary 5.1 hold as well as Assumption 3.3(ii), $b\in(1,\infty)$, and $c\in[1,2]$. Then
$\lim_{\tau\rightarrow 0}\lim\inf_{n\rightarrow\infty}\inf_{\hat{\gamma}}\sup_{p\in\mathcal{P}(b,c)}\mathbb{P}_{(W_{i})\sim p^{n}}\{\mathcal{R}(\hat{\gamma})>\tau\cdot n^{-\frac{bc}{bc+1}}\}=1.$
Corollary 5.2 (Optimality).
When $b\in(1,\infty)$ and $c\in(1,2]$, the upper bound in Theorem 5.1 matches the lower bound in Lemma 5.1. When $b\in(1,\infty)$ and $c=1$, they match up to a log factor.
@@ 5.2 Semiparametric inference beyond the RKHS
To showcase the role of Theorem 5.1 in semiparametric inference, I quote a result from the targeted and debiased machine learning literature. Unlike many previous works on kernel ridge balancing weights, inference allows $\gamma_{0}\not\in H$. It also allows for sample splitting, which is helpful when the regression estimator $\hat{\gamma}$ is not kernel ridge regression, and is instead estimated in a function space with high entropy. Theorem 5.1 is directly compatible with the quoted result, unlike previous guarantees for kernel ridge balancing weights. In this sense, KRRR provides debiased kernel methods.
@@
Definition 5.1 (Debiased machine learning).
@@
    Construct a $(1-a)100$% confidence interval as $\hat{\theta}\pm c_{a}\hat{\sigma}n^{-1/2}$, where $c_{a}$ is the $21-a/$ 2 quantile of the standard Gaussian.
Let $\gamma_{\mathcal{G}}$ be the best approximation to $\gamma_{0}$ in some space $\mathcal{G}$, which may not be an RKHS. Similarly, let $\alpha_{\mathcal{A}}$ be the best approximation to $\alpha_{0}$ in some space $\mathcal{A}$. Slightly abusing notation, I now let $\mathcal{R}(\hat{\gamma})=\mathbb{E}[\{\hat{\gamma}(W)-\gamma_{\mathcal{G}}(W)\}^{2}|(W_{i})_{i\in[n]}]$ and $\mathcal{R}(\hat{\alpha})=\mathbb{E}[\{\hat{\alpha}(W)-\alpha_{\mathcal{A}}(W)\}^{2}|(W_{i})_{i\in[n]}]$. [^8] In Theorem 5.1 above, $\mathcal{A}=H$. In Corollary 5.1 above, $\mathcal{G}=H$.
@@
Lemma 5.2 (Inference; Corollary 1 of [Chernozhukov et al., 2023]).
Suppose Assumption 3.1 holds. Suppose the following regularity conditions hold:
$\mathbb{E}[\{Y-\gamma_{0}(W)\}^{2}\mid W]\leq\bar{\sigma}^{2},\quad\|\alpha_{0}\|_{\infty}\leq\bar{\alpha},\quad\|\hat{\alpha}\|_{\infty}\leq\bar{\alpha}^{\prime},\quad\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left(\zeta/\sigma}\right)^{3}+\chi^{2}}\right\}n^{-1/2}\rightarrow 0.$
Suppose the following learning rate conditions hold:
@@
    $\mathopen{}\mathclose{{}\left(\bar{M}^{1/2}+\bar{\alpha}/\sigma+\bar{\alpha}^{\prime}}\right)\{\mathcal{R}(\hat{\gamma})+\|\gamma_{\mathcal{G}}-\gamma_{0}\|_{2}^{2}\}^{1/2}=o_{p}(1)$;
@@
    $\bar{\sigma}\{\mathcal{R}(\hat{\alpha})+\|\alpha_{\mathcal{A}}-\alpha_{0}\|_{2}^{2}\}^{1/2}=o_{p}(1)$;
  3. 3.
@@
Remark 5.3 (Mis-specification of features).
@@
To use Lemma 5.2 with classical kernel ridge balancing weights, allowing $\gamma_{0}\not\in H$, we require a rate $\mathcal{R}(\hat{\alpha})$, which is the main result of this paper (Theorem 5.1).
As written, Lemma 5.2 applies to pointwise approximations of nonparametric causal functions such as heterogeneous treatment effects (Example 3.2). In such case, $\sigma=\sigma_{h}\asymp h^{-1/2}$, so the rate of Gaussian approximation is $(nh)^{-1/2}$. This nonparametric rate is slower than the $n^{-1/2}$ obtained for regular cases such as ATE, similar to results of e.g. [van der Laan, 2006, van der Laan et al., 2018, Semenova and Chernozhukov, 2021, Nie and Wager, 2021, Foster and Syrgkanis, 2023, Kennedy, 2023] and many references therein, which could have been quoted instead. The point of Lemma 5.2 is not to provide new semiparametric theory but to demonstrate how the main result of this paper (Theorem 5.1) makes general semiparametric guarantees immediate. See Appendix A for more details on nonparametric causal functions. [^9] When appealing to Theorem 5.1 in the regular case, the rate for $\mathcal{R}(\hat{\alpha})$ is $r_{n}$. When appealing to Theorem 5.1 for causal functions, the rate for $\mathcal{R}(\hat{\alpha}_{h})$ is $h^{-2}r_{n}$.
Remark 5.4 (Rate-optimal regularization).
Lemma 5.2 does not require undersmoothing of the regularization in $\hat{\gamma}$ or $\hat{\alpha}$ in order to prove inference for $\hat{\theta}$, departing from e.g. [Hirshberg et al., 2019, Mou et al., 2023]. Instead, it requires $L_{2}$ rates for $\hat{\gamma}$ and $\hat{\alpha}$, the latter of which Theorem 5.1 provides. These rates are obtained with rate-optimal regularization when optimality is known, as formalized in Corollary 5.2. When $b<\infty$ and $c=1$, optimal regularization (up to log factors) is $\lambda=n^{-\frac{b}{b+1}}$.
@@
See [Bruns-Smith et al., 2023, Section 7] for a comprehensive and insightful discussion when $\hat{\gamma}=\tilde{\gamma}$ and $\hat{\alpha}=\tilde{\alpha}$.
@@
Proposition 5.1 (Consistency under mis-specification).
Suppose Assumption 3.1 holds. Suppose the following regularity conditions hold:
$\mathbb{E}[\{Y-\gamma_{\mathcal{G}}(W)\}^{2}\mid W]\leq\bar{\sigma}^{2},\quad\|\hat{\alpha}\|_{\infty}\leq\bar{\alpha}^{\prime},\quad\sigma_{\textsc{mis}}n^{-1/2}\rightarrow 0$
Suppose the following learning rate conditions hold: $\mathopen{}\mathclose{{}\left(\bar{M}^{1/2}+\bar{\alpha}^{\prime}}\right)\mathcal{R}(\hat{\gamma})^{1/2}=o_{p}(1)$ and $\bar{\sigma}\mathcal{R}(\hat{\alpha})^{1/2}=o_{p}(1)$. If either $\gamma_{\mathcal{G}}=\gamma_{0}$ or $\alpha_{\mathcal{A}}=\alpha_{0}$, then $\hat{\theta}=\theta_{0}+o_{p}(1)$.
See Appendix C for the proof and the exact nonasymptotic rate of convergence, which depends on $\mathcal{R}(\hat{\gamma})$ and $\mathcal{R}(\hat{\alpha})$. Previous works on kernel ridge balancing weights typically require $\gamma_{0}\in H$. By contrast, Proposition 5.1 tolerates nonvanishing $\|\gamma_{\mathcal{G}}-\gamma_{0}\|_{2}$ or nonvanishing $\|\alpha_{\mathcal{A}}-\alpha_{0}\|_{2}$, where neither $\mathcal{G}$ nor $\mathcal{A}$ may be $H$. To achieve this robustness to mis-specification for KRRR, we require a rate $\mathcal{R}(\hat{\alpha})$, which is the main result of this paper.
Again, as written, Proposition 5.1 applies to pointwise approximations of nonparametric causal functions. See Appendix A for details.
Proposition 5.1 summarizes the nonasymptotic Proposition C.1, which is a modest refinement of the celebrated double robustness guarantee. It is merely for illustrative purposes; stronger asymptotic results are available in the literature. For inference under mis-specification, see e.g. [van der Laan, 2014, Benkeser et al., 2017, Dukes et al., 2021].
6 Proof of main result via the counterfactual effective dimension

### v4

@@ 4.3   Standalone closed form solution
In summary, the standalone closed form for KRRR is not necessary for causal estimation in the special case of Corollary 5. However, Corollary 6 shows that if $\hat{\gamma}\neq\tilde{\gamma}$, or if $\hat{\gamma}$ or $\tilde{\alpha}$ is estimated on a held out sample, then a standalone closed form for KRRR that can be evaluated away from the training data is necessary for causal estimation. I derive it below using techniques of [Chernozhukov et al., 2020], as a secondary contribution.
Even for the special case of Corollary 5, this paper’s primary contribution is the population $L_{2}$ rate $\|\tilde{\alpha}-\alpha_{H}\|_{2}$ for kernel ridge balancing weights (and more generally KRRR) in Section 5. Generalization error is essential for appealing to strong semiparametric guarantees that allow some mis-specification and justify inference on causal functions.
I use some additional notation for features. Define the operator $\Phi:H\rightarrow\mathbb{R}^{n}$ with $i$th component $\langle\phi(W_{i}),\cdot\rangle_{H}$ and likewise the operator $\Phi^{(m)}:H\rightarrow\mathbb{R}^{n}$ with $i$th component $\langle\phi^{(m)}(W_{i}),\cdot\rangle_{H}$. Concatenate $\Phi$ and $\Phi^{(m)}$ as $\Psi:H\rightarrow\mathbb{R}^{2n}$, with adjoint $\Psi^{*}:\mathbb{R}^{2n}\rightarrow H$.
Lemma 5 (Standalone closed form exists; c.f. Lemma 2 of [Chernozhukov et al., 2020]).
Suppose the conditions of Lemma 3 hold. There exists some $\rho\in\mathbb{R}^{2n}$ such that $\tilde{\alpha}=\Psi^{*}\rho$.
@@
The familiar kernel matrix $K^{(1)}=\Phi\Phi^{*}\in\mathbb{R}^{n\times n}$ encodes inner products between actual observations. The extended kernel matrix $K=\Psi\Psi^{*}\in\mathbb{R}^{2n\times 2n}$ encodes inner products between actual observations and counterfactual observations. Each entry of $K$ is computed from the kernel $k$ and formula $m$ as discussed in Section 3.3. For intuition, write
@@
Proposition 2 (Standalone closed form for KRRR).
Suppose the conditions of Lemma 3 hold. Given observations $(W_{i})$ and evaluation location $w$, as well as the formula $m$, kernel $k$, and regularization $\lambda$, the KRRR closed form is as follows.
@@
    Calculate $\Omega^{2n\times 2n}$ and $u(w),v\in\mathbb{R}^{2n}$ by
@@
The standalone closed form solution for kernel ridge balancing weights, and KRRR more generally, appears to have been previously unknown. Proposition 2 demonstrates that it can be easily computed using only the formula $m$ and kernel $k$, without directly evaluating the feature maps $\phi(w)$ or $\phi^{(m)}(w)$. In this sense, it is a practical contribution for inference. See Appendix E for its specialization to Example 2.
5   Generalization error with mis-specification
@@ 5.1   Main result
I now provide the main result: population $L_{2}$ rates for KRRR that coincide with the optimal rate, where optimal rates are known. I study the class of Riesz representers defined by Assumption 1, imposing familiar smoothness and spectral decay conditions for the RKHS setting defined by Assumptions 2 and 3. I also impose mild regularity conditions that help in the algorithm derivation, defined by Assumptions 4 and 5.
@@
Theorem 2 (Upper bound for KRRR).
If Assumptions 123(i), 4, and 5 hold with $\bar{M}$ bounded above, then $\lim_{\tau\rightarrow\infty}\lim\sup_{n\rightarrow\infty}\sup_{p\in\mathcal{P}(b,c)}\mathbb{P}_{(W_{i})\sim p^{n}}\{\mathcal{R}(\tilde{\alpha})>\tau\cdot r_{n}\}=0$, where
$r_{n}=\begin{cases}n^{-1}&\text{ if }\quad b=\infty\text{ and }\lambda=n^{-\frac{1}{2}};\\ n^{-\frac{bc}{bc+1}}&\text{ if }\quad b\in(1,\infty),\;c\in(1,2]\text{ and }\lambda=n^{-\frac{b}{bc+1}};\\ \ln^{\frac{b}{b+1}}(n)\cdot n^{-\frac{b}{b+1}}&\text{ if }\quad b\in(1,\infty),\;c=1\text{ and }\lambda=\ln^{\frac{b}{b+1}}(n)\cdot n^{-\frac{b}{b+1}}.\end{cases}$
Remark 1 (Causal functions).
When the causal parameter has a formula of the form $m_{h}(w,f)=\ell_{h}(w)\check{m}(w,f)$, as in Example 2, $\bar{M}_{h}\uparrow\infty$. In such case, I use the estimator $\hat{\alpha}_{h}(W)=\ell_{h}(W)\tilde{\alpha}(W)$, where $\tilde{\alpha}$ is KRRR for the formula $\check{m}(w,f)$. When $\tilde{\alpha}$ has the rate $r_{n}$, $\hat{\alpha}_{h}$ has the rate $h^{-2}r_{n}$ under weak regularity conditions. See Appendix A for details.
@@
Theorem 2 does not require Assumption 3(ii). Moreover, as an intermediate step, I prove a nonasymptotic bound (Proposition 5) without imposing Assumptions 2 and 3(i).
When the RKHS is finite dimensional, i.e. $b=\infty$ in Assumption 3, KRRR achieves the parametric rate $r_{n}=n^{-1}$. When the RKHS is infinite dimensional with at least polynomial spectral decay, i.e. $b<\infty$ in Assumption 3, KRRR achieves a familiar rate $r_{n}=n^{-\frac{bc}{bc+1}}$. This rate depends on $c$ in Assumption 2, which quantifies the smoothness of the best kernel approximation $\alpha_{H}\in H$ to the Riesz representer $\alpha_{0}\in L_{2}$. If no additional smoothness of $\alpha_{H}$ is known, i.e. all we have is that $\alpha_{H}\in H$ by construction, then $c=1$ and the rate has an extra logarithmic factor.
The rate $r_{n}=n^{-\frac{bc}{bc+1}}$ generalizes the familiar rate from Sobolev analysis. For example, take $\mathcal{W}=\mathbb{R}^{p}$ and denote by $\mathbb{H}_{2}^{s}$ the Sobolev space with $s>p/2$ square integrable derivatives, as before. Suppose that KRRR $\tilde{\alpha}$ is estimated using the RKHS $H=\mathbb{H}_{2}^{s}$. Suppose that $\alpha_{H}\in\mathbb{H}_{2}^{s_{0}}$, i.e. the best kernel approximation to the Riesz representer $\alpha_{0}\in L_{2}$ has $s_{0}>s$ square integrable derivatives. The population $L_{2}$ rate for $\tilde{\alpha}$ is then $r_{n}=n^{-\frac{2s_{0}}{2s_{0}+p}}$. When $s_{0}=s$, which is a lower bound on $s_{0}$ by construction, $r_{n}=\ln(n)^{\frac{2s}{2s+p}}\cdot n^{-\frac{2s}{2s+p}}$. Theorem 2 applies to settings beyond Sobolev spaces over Euclidean domains.
Corollary 7 (Special case: Kernel ridge regression; Theorem 1 of [Caponnetto and De Vito, 2007]).
Suppose the conditions of Theorem 2 hold, replacing Assumption 1 with $|Y|\leq\bar{Y}$ almost surely. Then $\lim_{\tau\rightarrow\infty}\lim\sup_{n\rightarrow\infty}\sup_{p\in\mathcal{P}(b,c)}\mathbb{P}_{(W_{i})\sim p^{n}}\{\mathcal{R}(\tilde{\gamma})>\tau\cdot r_{n}\}=0.$
Remark 2 (Excess risk).
@@
Just as KRRR generalizes kernel ridge regression, Theorem 2 generalizes [Caponnetto and De Vito, 2007, Theorem 1]. The contribution is meaningful because KRRR also generalizes the classical kernel ridge balancing weights, for which generalization error was not fully articulated. In settings where lower bounds are known, the upper bounds in Theorem 2 often achieve them.
Lemma 6 (Lower bound; Theorem 2 of [Caponnetto and De Vito, 2007]).
Suppose the conditions of Corollary 7 hold as well as Assumption 3(ii), $b\in(1,\infty)$, and $c\in[1,2]$. Then $\lim_{\tau\rightarrow 0}\lim\inf_{n\rightarrow\infty}\inf_{\hat{\gamma}}\sup_{p\in\mathcal{P}(b,c)}\mathbb{P}_{(W_{i})\sim p^{n}}\{\mathcal{R}(\hat{\gamma})>\tau\cdot n^{-\frac{bc}{bc+1}}\}=1.$
Corollary 8 (Optimality).
When $b\in(1,\infty)$ and $c\in(1,2]$, Theorem 2 matches Lemma 6. When $b\in(1,\infty)$ and $c=1$, they match up to a log factor.
@@ 5.2   Semiparametric inference beyond the RKHS
To showcase the role of Theorem 2 in semiparametric inference, I quote a result from the targeted and debiased machine learning literature. Unlike many previous works on kernel ridge balancing weights, inference allows $\gamma_{0}\not\in H$. It also allows for sample splitting, which is helpful when the regression estimator $\hat{\gamma}$ is not kernel ridge regression, and is instead estimated in a function space with high entropy. Theorem 2 is directly compatible with the quoted result, unlike previous guarantees for kernel ridge balancing weights. In this sense, KRRR provides debiased kernel methods.
@@
Definition 4 (Debiased machine learning).
@@
    Construct the $(1-a)100$% confidence interval as $\hat{\theta}\pm c_{a}\hat{\sigma}n^{-1/2}$, where $c_{a}$ is the $21-a/$ 2 quantile of the standard Gaussian.
Let $\gamma_{\mathcal{G}}$ be the best approximation to $\gamma_{0}$ in some space $\mathcal{G}$, which may not be an RKHS. Similarly, let $\alpha_{\mathcal{A}}$ be the best approximation to $\alpha_{0}$ in some space $\mathcal{A}$. Slightly abusing notation, I now let $\mathcal{R}(\hat{\gamma})=\mathbb{E}[\{\hat{\gamma}(W)-\gamma_{\mathcal{G}}(W)\}^{2}|(W_{i})_{i\in[n]}]$ and $\mathcal{R}(\hat{\alpha})=\mathbb{E}[\{\hat{\alpha}(W)-\alpha_{\mathcal{A}}(W)\}^{2}|(W_{i})_{i\in[n]}]$. [^9] In Theorem 2 above, $\mathcal{A}=H$. In Corollary 7 above, $\mathcal{G}=H$.
@@
Lemma 7 (Inference; Corollary 1 of [Chernozhukov et al., 2023]).
Suppose Assumption 1 holds. Suppose the following regularity conditions: $\mathbb{E}[\{Y-\gamma_{0}(W)\}^{2}\mid W]\leq\bar{\sigma}^{2}$, $\|\alpha_{0}\|_{\infty}\leq\bar{\alpha}$, $\|\hat{\alpha}\|_{\infty}\leq\bar{\alpha}^{\prime}$, and $\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left(\zeta/\sigma}\right)^{3}+\chi^{2}}\right\}n^{-1/2}\rightarrow 0.$ Suppose the following learning rate conditions:
@@
    $\mathopen{}\mathclose{{}\left(\bar{M}^{1/2}+\bar{\alpha}/\sigma+\bar{\alpha}^{\prime}}\right)\{\mathcal{R}(\hat{\gamma})+\|\gamma_{\mathcal{G}}-\gamma_{0}\|_{2}^{2}\}^{1/2}=o_{p}(1)$; $\bar{\sigma}\{\mathcal{R}(\hat{\alpha})+\|\alpha_{\mathcal{A}}-\alpha_{0}\|_{2}^{2}\}^{1/2}=o_{p}(1)$;
@@
Remark 3 (Mis-specification of features).
@@
To use Lemma 7 with classical kernel ridge balancing weights, allowing $\gamma_{0}\not\in H$, we require a rate $\mathcal{R}(\hat{\alpha})$, which is the main result of this paper (Theorem 2).
As written, Lemma 7 applies to pointwise approximations of nonparametric causal functions such as heterogeneous treatment effects (Example 2). In such case, $\sigma=\sigma_{h}\asymp h^{-1/2}$, so the rate of Gaussian approximation is $(nh)^{-1/2}$. In practice, I take $h\asymp n^{-1/5}$ so the rate becomes $n^{-2/5}$; see Section 7. This nonparametric rate is slower than the $n^{-1/2}$ obtained for regular cases such as ATE, similar to results of e.g. [van der Laan et al., 2018, Nie and Wager, 2021, Foster and Syrgkanis, 2023, Kennedy, 2023] and many references therein, which could have been quoted instead. The point of Lemma 7 is not to provide new semiparametric theory but to demonstrate how the main result of this paper (Theorem 2) makes general semiparametric guarantees immediate. See Appendix A for more details on nonparametric causal functions. [^10] When appealing to Theorem 2 in the regular case, the rate for $\mathcal{R}(\hat{\alpha})$ is $r_{n}$. When appealing to Theorem 2 for causal functions, the rate for $\mathcal{R}(\hat{\alpha}_{h})$ is $h^{-2}r_{n}$.
Remark 4 (Rate-optimal ridge regularization).
Lemma 7 does not require undersmoothing of the regularization in $\hat{\gamma}$ or $\hat{\alpha}$ in order to prove inference for $\hat{\theta}$, departing from e.g. [Hirshberg et al., 2019, Mou et al., 2023]. Instead, it requires $L_{2}$ rates for $\hat{\gamma}$ and $\hat{\alpha}$, the latter of which Theorem 2 provides. These rates are obtained with rate-optimal ridge regularization when optimality is known, as formalized in Corollary 8. When $b<\infty$ and $c=1$, optimal ridge regularization (up to log factors) is $\lambda=n^{-\frac{b}{b+1}}$.
@@
Proposition 3 (Consistency under mis-specification).
Suppose Assumption 1 holds. Suppose the following regularity conditions: $\mathbb{E}[\{Y-\gamma_{\mathcal{G}}(W)\}^{2}\mid W]\leq\bar{\sigma}^{2}$, $\|\hat{\alpha}\|_{\infty}\leq\bar{\alpha}^{\prime}$, and $\sigma_{\textsc{mis}}n^{-1/2}\rightarrow 0$. Suppose the following learning rate conditions: $\mathopen{}\mathclose{{}\left(\bar{M}^{1/2}+\bar{\alpha}^{\prime}}\right)\mathcal{R}(\hat{\gamma})^{1/2}=o_{p}(1)$ and $\bar{\sigma}\mathcal{R}(\hat{\alpha})^{1/2}=o_{p}(1)$. If either $\gamma_{\mathcal{G}}=\gamma_{0}$ or $\alpha_{\mathcal{A}}=\alpha_{0}$, then $\hat{\theta}=\theta_{0}+o_{p}(1)$.
See Appendix C for the proof and the exact nonasymptotic rate of convergence, which depends on $\mathcal{R}(\hat{\gamma})$ and $\mathcal{R}(\hat{\alpha})$. Previous works on kernel ridge balancing weights typically require $\gamma_{0}\in H$. By contrast, Proposition 3 tolerates nonvanishing $\|\gamma_{\mathcal{G}}-\gamma_{0}\|_{2}$ or nonvanishing $\|\alpha_{\mathcal{A}}-\alpha_{0}\|_{2}$, where neither $\mathcal{G}$ nor $\mathcal{A}$ may be $H$. To achieve this robustness to mis-specification for KRRR, we require a rate $\mathcal{R}(\hat{\alpha})$, i.e. the main result of this paper.
Again, as written, Proposition 3 applies to pointwise approximations of nonparametric causal functions. See Appendix A for details.
Proposition 3 summarizes the nonasymptotic Proposition 6, which is a modest refinement of the celebrated double robustness guarantee. It is merely for illustrative purposes; stronger asymptotic results are available in the literature. For inference under mis-specification, see e.g. [van der Laan, 2014, Benkeser et al., 2017, Dukes et al., 2021].
6   Proof via the counterfactual effective dimension